OpenStudy (anonymous):

What is the value of x ? A. 77 B. 70 C. 103 D. 84

OpenStudy (anonymous):

Here is the problem

Nnesha (nnesha):

all triangles add up to what degree ??

OpenStudy (anonymous):

103?

Nnesha (nnesha):

i'm not asking from that answer choices basic information about triangles in triangles all 3 angles equal to what ??

Nnesha (nnesha):

do you know that ?^

OpenStudy (anonymous):

Always add up to 180

Nnesha (nnesha):

yep right :)

Nnesha (nnesha):

so add  \[\huge\rm 42+32+x = 180\] so for x

Nnesha (nnesha):

add both angles and then subtract from 180

OpenStudy (anonymous):

So 42 + 32 = 74 so do I subtract with 180

Nnesha (nnesha):

not 32 hmm *facepalm* 35 sorry

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\large\rm 42+32+x = 180\] so for x \(\color{blue}{\text{End of Quote}}\) its 35 \[\huge\rm 42+35+x = 180\]

Nnesha (nnesha):

so yes 42 + 35 and then subtract 180

OpenStudy (anonymous):

42 + 35 = 77 and 180 - 77 = 103

Nnesha (nnesha):

yep right wait did you tell m final answer before ?? :O

OpenStudy (anonymous):

Yeah

Nnesha (nnesha):

hmm sorry i was asking diff question so i thought u were saying all triangles add up to 103 hmm so sorry

OpenStudy (anonymous):

It's okay you still helped me thanks

Nnesha (nnesha):

thanks and np :)